### v2

@@ 2.3.3 Phenotyping and Medical Coding
The phenotyping task primarily involves identifying phenotypic abnormalities from a patient’s various medical records, which aids in the identification of rare diseases. There exists the Human Phenotype Ontology (HPO) project [^1] https://hpo.jax.org that systematically categorizes human phenotypes with detailed annotations. Thus, the phenotyping task can be framed as a multi-label classification task where we need to find mapping $f:\mathbf{X}\rightarrow 2^{\mathbf{Y}}$, with $\mathbf{X}$ representing the set of patient medical records, where each medical record $x_{i}\in\mathbf{X}$ corresponds to the medical data (unstructured text in most cases) from the $i$-th record, $\mathbf{Y}$ denotes the set of all possible phenotype labels from HPO, where each label $y_{j}\in\mathbf{Y}$ represents a specific phenotype. The objective of the phenotyping task is to map each medical record $x_{i}$ to a subset of phenotype labels $Y_{i}\subseteq\mathbf{Y}$, such that $Y_{i}$ represents the set of phenotypes exhibited by the patient associated with $x_{i}$:
$Y_{i}=f(x_{i})$
@@
Medical coding is another multi-label classification task that involves identifying a set of International Classification of Diseases (ICD) codes [^2] https://icd.who.int associated with a medical record. This task can be formulated similarly to phenotyping, with the key difference being that $\mathbf{Y}$ represents a set of ICD codes rather than phenotype labels. Besides observing trends similar to those in the phenotyping subdomain, it is also noteworthy that there is a shift towards explainable medical coding, as highlighted in [132].
@@ 2.4 Question Answering and Information Extraction
Masked models are well-suited for question answering tasks. They can be fine-tuned on specific QA datasets, where the input is a concatenation of the question and context (a paragraph or document containing the answer). The model is then trained to identify the span of text that answers the question, typically by adding a start and end token classifier to the output embeddings of the MLM. These classifiers predict the beginning and end positions of the answer in the text. On the other hand, generative models leverage their extensive pre-training on diverse data. By inputting a question (along with the document of interest when necessary) and following it with a prompt that encourages the model to generate an answer, these models can produce responses without needing explicit pointers to answer spans.
In the medical domain, IE and QA systems are instrumental for extracting data from electronic health records, such as medication lists and diagnostic details, essential for patient management and treatment planning. A notable example is quEHRy, a QA system designed to query EHRs using natural language. The primary goal of quEHRy is to provide precise and interpretable answers to clinicians’ questions from structured EHR data [150]. Beyond the successful applications of BERT-based models like BioBERT, BiomedBERT, and PubMedBERT for QA and IE, generative models also show proficiency. Agrawal et al. demonstrated the effectiveness of generative LLMs such as InstructGPT and GPT-3 in zero-shot and few-shot information extraction from clinical texts. When tested on the re-annotated CASI dataset, these models showed considerable potential in tasks requiring structured outputs [185]. Furthermore, Ge et al. compared the effectiveness of LLMs versus manual chart reviews for extracting data elements from EHRs, focusing specifically on hepatocellular carcinoma imaging reports. Using the GPT-3.5-turbo model, implemented as “Versa Chat” within a secure UCSF  [^3] https://www.ucsf.edu/ environment to protect patient health information, the study analyzed 182 CT or MRI abdominal imaging reports from the Functional Assessment in Liver Transplantation study. It extracted six distinct data elements, including the maximum LI-RADS [^4] https://www.acr.org/Clinical-Resources/Reporting-and-Data-Systems/LI-RADS score, number of hepatocellular carcinoma lesions, and presence of macrovascular invasion. The performance was evaluated by calculating accuracy, precision, recall, and F1 scores, showing high overall accuracy (0.889) with variations depending on the complexity of the data elements [186].
@@ 2.5 Summarization and Paraphrasing
Masked language models are proficient at extractive summarization. They evaluate sentences within a text to determine their relevance and informativeness. By scoring each sentence, these models identify and concatenate the most important sentences to form a coherent summary. In contrast, abstractive summarization and paraphrasing typically employ generative (decoder-based) or sequence-to-sequence (encoder-decoder) models. These models are trained to understand the entire narrative or document and then recreate its essence in a different form.
@@
Summarization and paraphrasing in the medical domain are largely driven by advancements in these tasks in general. Devaraj et al. introduce a new dataset derived from the Cochrane Database of Systematic Reviews, featuring pairs of technical abstracts and plain language summaries. They propose a novel metric based on masked language models to better distinguish between technical and simplified texts. The study utilizes baseline encoder-decoder Transformer models for text simplification and introduces an innovative approach to penalize the generation of jargon terms. The code and data are publicly available for further research [153].
@@
The paper titled “Biomedical Text Readability After Hypernym Substitution with Fine-Tuned Large Language Models” investigates simplifying biomedical text using LLMs to enhance patient understanding. The authors fine-tuned three LLM variants to replace complex biomedical terms with their hypernyms. The models used include GPT-J-6b, SciFive T5, and an approach combining sequence-to-sequence and sciBERT models. The study processed 1,000 biomedical definitions from the Unified Medical Language System and evaluated readability improvements using metrics such as the Flesch-Kincaid Reading Ease and Grade Level, Automated Readability Index, and Gunning Fog Index. Results showed substantial readability improvements, with the GPT-J-6b model performing best in reducing sentence complexity [156].
@@
Another interesting application of paraphrasing is the anonymization of medical documents, which is crucial for balancing ethical principles and research needs. Wiest et al. present an approach to de-identify medical free text using LLMs. The authors benchmarked eight locally deployable LLMs, including Llama-3 8B, Llama-3 70B, Llama-2 7B, Llama-2 70B, and Mistral 7B, on a dataset of 100 clinical letters from a German hospital. They developed the LLM-Anonymizer pipeline, which achieved a success rate of 98.05% in removing personal identifying information using Llama-3 70B. The tool is open-source, operates on local hardware, and does not require programming skills, making it accessible and practical for use in medical institutions. The study demonstrates the potential of LLMs to effectively de-identify medical texts, outperforming traditional NLP methods and providing a robust solution for privacy-preserving data sharing in healthcare [157].
@@
Despite advancements in summarization and paraphrasing, some challenges persist, particularly in preserving factual accuracy and precision. Jeblick et al. explored the effectiveness of using ChatGPT (version December 15th, 2022) to simplify radiology reports into language understandable by non-experts. A radiologist created three hypothetical radiology reports, which were then simplified by prompting ChatGPT. Fifteen radiologists evaluated the quality of these simplified reports based on criteria such as factual correctness, completeness, and potential harm to patients. The study used Likert scale analysis and inductive free-text categorization to assess the simplified reports. Overall, the radiologists found the simplified reports to be factually correct and complete, with minimal potential for harm. However, some issues were noted, including incorrect information, omissions of relevant medical data, and occasionally misleading or vague statements. These issues highlight the need for careful supervision by medical professionals when using language models to simplify complex medical texts [155]. A recent study by Landman et al. discusses a challenge organized by Pfizer to explore the use of LLMs for automating the summarization of safety tables in clinical study reports. Various teams employed GPT models with prompt engineering techniques to generate summary texts. The datasets included safety outputs from 72 reports from recent clinical studies, split into 70% for training and 30% for testing. The study concluded that while LLMs show promise in automating the summarization of clinical study report tables, human involvement and further research are necessary to optimize their application [151].
@@ 2.6 Conversation
The task of conversation, or dialogue generation, can be formulated as follows. Given a dialogue history $H=(h_{1},h_{2},\ldots,h_{n})$, where each $h_{i}$ represents an utterance in the conversation, the goal is to generate a response $R$.
Typically, pre-trained decoder-based large language models are fine-tuned using specialized datasets to develop their conversational capabilities. In the medical domain, conversational applications facilitate interactive communication with patients. For example, conversational AI can be deployed as virtual health assistants that provide initial consultations based on symptoms described by patients. These systems can ask relevant follow-up questions, assess symptoms, and offer preliminary advice or direct patients to seek professional care when necessary. Additionally, these conversational tools can be utilized for patient education, explaining complex medical conditions and treatments in simple language to enhance understanding and compliance. Another notable application is in mental health support, where conversational AI can offer coping strategies and basic support, thereby augmenting traditional therapy sessions.
@@ 2.6.1 Chatbots and Health Assistants
Cung et al. assessed the performance of three commercial systems - ChatGPT, Bing, and Bard in the context of skeletal biology and disorders. The study involved posing 30 questions across three categories, with the responses graded for accuracy by four reviewers. While ChatGPT 4.0 had the highest overall median score, the study revealed that the quality and relevance of responses from all three chatbots varied widely, presenting issues such as inconsistency and failure to account for patient demographics [163]. Another study explored using ChatGPT for patient-provider communication. A survey of 430 participants found that ChatGPT responses were often indistinguishable from those of healthcare providers, indicating a level of trust in chatbots for answering lower-risk health questions [164].
@@
Another research direction in the field of medical chatbots is the implementation of conversational agents specifically designed for the medical domain. Abbasian et al. proposed a complex LLM-based multimodal framework for CHAs, concentrating on critical thinking, knowledge acquisition, and multi-step problem-solving. This framework aims to enable CHAs to provide personalized healthcare responses and handle intricate tasks such as stress level estimation [113].
Domain-specific LLMs, such as the ChatDoctor model, can also serve as chatbots. This model integrates a self-directed information retrieval mechanism, allowing it to access up-to-date information from online and curated offline medical databases. Evaluated using BERTScore[188], ChatDoctor exhibited a higher F1 score compared to ChatGPT-3.5, demonstrating the effectiveness of smaller domain-specific models as alternatives to large commercial solutions [94].
@@ 2.6.2 Mental Health Bots
The idea of using a machine as a personal psychologist dates back to at least the 1960s when Weizenbaum proposed a simple rule-based system called ELIZA [190]. Contemporary advancements in mental health chatbots are largely driven by LLMs. Yang et al. investigated the capabilities of current LLMs in automated mental health analysis. Their study involved evaluating LLMs across diverse datasets for tasks such as emotional reasoning and detecting mental health conditions, employing various similarity metrics including BLEU, ROUGE family, BERTScore derivatives, BART-score [191], and human assessments. They discovered that while ChatGPT displays robust in-context learning abilities, it still encounters challenges in emotion-related tasks and requires careful prompt engineering to improve performance to enhance its performance [161].
@@
Saha et al. introduced a Virtual Assistant for supporting individuals with Major Depressive Disorder, using a dataset called MotiVAte [160]. Their system, based on modified GPT-2 model and reinforced learning, shows promising results in generating empathetic and motivational responses, as evidenced by both automated evaluations based on text similarity and human evaluations based on fluency, adaptability, and degree of motivation. Sharma et al. introduced a dataset for training a GPT-3-based model for generating reframes with controlled linguistic attributes [192]. Deployed on the Mental Health America website [^5] https://mhanational.org/, this allowed for a randomized field study to gather findings on human preferences. Another team explored the fine-tuning of open-source LLMs on psychotherapy assistant instructions, using a dataset from Alexander Street Press [^6] https://alexanderstreet.com/ therapy and counseling sessions. Their results indicated that LLMs fine-tuned on domain-specific instructions surpassed their non-fine-tuned counterparts in psychotherapy tasks, underscoring the significance of professional and context-specific training for these models [162].
@@
Promising outcomes have been observed through collaborations between humans and AI. A recent study conducted a randomized controlled trial involving human peer supporters, demonstrating that an AI-in-the-loop agent led to a 19.60% increase in conversational empathy in interactions between individuals seeking mental health support and support specialists. This was achieved by providing suggestions for response improvements to peer supporters [193]. This research reveals that human-AI collaboration is a crucial area for potential exploration, particularly in the medical domain.
@@
The evolving reasoning capabilities of LLMs have sparked interest in their use for disease diagnostics. Levine et al. conducted experiments with the GPT-3 model to assess its diagnostic and triage accuracy [165]. Their results indicate that GPT-3’s diagnostic accuracy is comparable to that of physicians but lags in triage accuracy. GPT-3 correctly identified the diagnosis in its top three choices for 88% of the cases, surpassing non-experts (54%) but slightly underperforming compared to professional physicians (96%). In triage performance, GPT-3 achieved an accuracy of 70%, on par with non-experts (74%) but significantly lower than physicians (91%). Despite GPT-3’s notable performance, the study raises ethical concerns, particularly regarding the model’s potential to perpetuate existing data biases, exhibiting racial and gender biases and occasionally producing misleading or incorrect information.
@@
A recent study by Liu et al. [194] introduced a framework named PharmacyGPT, which leverages the current GPT family models to emulate the role of clinical pharmacists. This research utilized real data from the ICU at the University of North Carolina Chapel Hill (UNC) Hospital. PharmacyGPT was applied to tackle various challenges in the realm of pharmacy, encompassing patient outcome studies, AI-based medication prescription generation, and interpretable patient clustering analysis. The study revealed that the GPT-4 model, when provided with dynamic context and similar samples, attained the highest accuracy among all models tested. However, the precision and recall scores were not notably high across the approaches. This outcome may be caused by the binary nature of mortality prediction, a significant imbalance in the dataset, and the complex, individualized nature of ICU pharmacy regimens. The research highlights the need for custom evaluation metrics to assess the performance of AI-generated medication plans, enhancing understanding of the models’ strengths and limitations [194].
3 Challenges and Opportunities

### v3

@@ 3.3.3 Phenotyping and Medical Coding
The phenotyping task primarily involves identifying phenotypic abnormalities from a patient’s various medical records, which aids in the identification of rare diseases. There exists the Human Phenotype Ontology (HPO) project [^1] https://hpo.jax.org that systematically categorizes human phenotypes with detailed annotations. The phenotyping task can be framed as a multi-label classification task, which extends the single-label classification task as follows. Given a sequence of tokens $X=(x_{1},x_{2},\ldots,x_{n})$, the task is to assign a subset of labels $Y\subseteq\mathcal{Y}$, where $\mathcal{Y}$ is the set of all possible labels.
For each label $k\in\mathcal{Y}$, the model predicts a probability $P(k\mid X)$ representing the likelihood of the sequence $X$ being associated with label $k$. The set of labels $Y$ is typically determined by applying a predefined threshold $\theta$ to these probabilities:
$Y=\{k\in\mathcal{Y}\mid P(k\mid X)\geq\theta\}.$
Thus, in the phenotyping task, the sequence $X$ represents a medical record of a patient, and $\mathcal{Y}$ denotes the set of all possible phenotype labels from the HPO. The objective is to identify a subset $Y\subseteq\mathcal{Y}$ of HPO labels associated with the medical record.
@@
Medical coding is another multi-label classification task that involves identifying a set of International Classification of Diseases (ICD) codes [^2] https://icd.who.int associated with a medical record. This task can be formulated similarly to phenotyping, with the key difference being that $\mathcal{Y}$ represents a set of ICD codes rather than HPO labels. Besides observing trends similar to those in the phenotyping subdomain, it is also noteworthy that there is a shift towards explainable medical coding, as highlighted in [154].
@@ 3.4 Question Answering and Information Extraction
Encoder-based models are well-suited for question answering tasks. They can be fine-tuned on specific QA datasets, where the input is a concatenation of the question and context (a paragraph or document containing the answer). The model is then trained to identify the span of text that answers the question, typically by adding a start and end token classifier to the output embeddings of the model. These classifiers predict the beginning and end positions of the answer in the text. On the other hand, generative models with a decoder component leverage their extensive pre-training on diverse data. By inputting a question (along with the document of interest when necessary) and following it with a prompt that encourages the model to generate an answer, these models can produce responses without needing explicit pointers to answer spans.
In the medical domain, IE and QA systems are instrumental for extracting data from electronic health records, such as medication lists and diagnostic details, essential for patient management and treatment planning. A notable example is quEHRy, a QA system designed to query EHRs using natural language. The primary goal of quEHRy is to provide precise and interpretable answers to clinicians’ questions from structured EHR data [172]. Beyond the successful applications of BERT-based models like BioBERT, BiomedBERT, and PubMedBERT for QA and IE, generative models also show proficiency. Agrawal et al. [213] demonstrated the effectiveness of generative LLMs such as InstructGPT and GPT-3 in zero-shot and few-shot information extraction from clinical texts, measured through accuracy and F1 score. When tested on the re-annotated CASI dataset, these models showed considerable potential in tasks requiring structured outputs. Furthermore, Ge et al. [214] compared the effectiveness of LLMs versus manual chart reviews for extracting data elements from EHRs, focusing specifically on hepatocellular carcinoma imaging reports. Using the GPT-3.5-turbo model, implemented as “Versa Chat” within a secure UCSF  [^3] https://www.ucsf.edu/ environment to protect patient health information, the study analyzed 182 CT or MRI abdominal imaging reports from the Functional Assessment in Liver Transplantation study. It extracted six distinct data elements, including the maximum LI-RADS [^4] https://www.acr.org/Clinical-Resources/Reporting-and-Data-Systems/LI-RADS score, number of hepatocellular carcinoma lesions, and presence of macrovascular invasion. The performance was evaluated by calculating accuracy, precision, recall, and F1 scores, showing high overall accuracy (0.889) with variations depending on the complexity of the data elements.
@@ 3.5 Summarization and Paraphrasing
Encoder-based models are proficient at extractive summarization. They evaluate sentences within a text to determine their relevance and informativeness. By scoring each sentence, these models identify and concatenate the most important sentences to form a coherent summary. In contrast, abstractive summarization and paraphrasing typically employ decoder-based or sequence-to-sequence (encoder-decoder) models. These models are trained to understand the entire narrative or document and then recreate its essence in a different form.
@@
Summarization and paraphrasing in the medical domain are largely driven by advancements in these tasks in general. Devaraj et al. [175] introduce a new dataset derived from the Cochrane Database of Systematic Reviews [^5] https://www.cochranelibrary.com/cdsr/about-cdsr, featuring pairs of technical abstracts and plain language summaries. They propose a novel metric based on encoder-based language models to better distinguish between technical and simplified texts. The study utilizes baseline encoder-decoder Transformer models for text simplification and introduces an innovative approach to penalize the generation of jargon terms. The code and data are publicly available for further research.
@@
The paper titled “Biomedical Text Readability After Hypernym Substitution with Fine-Tuned Large Language Models” investigates simplifying biomedical text using LLMs to enhance patient understanding. The authors fine-tuned three LLM variants to replace complex biomedical terms with their hypernyms. The models used include GPT-J-6b, SciFive T5 [215], and an approach combining sequence-to-sequence and sciBERT [216] models. The study processed 1,000 biomedical definitions from the Unified Medical Language System (UMLS) and evaluated readability improvements using metrics such as the Flesch-Kincaid Reading Ease and Grade Level, Automated Readability Index, and Gunning Fog Index. Results showed substantial readability improvements, with the GPT-J-6b model performing best in reducing sentence complexity [178].
@@
Another interesting application of paraphrasing is the anonymization of medical documents, which is crucial for balancing ethical principles and research needs. Wiest et al.[179] present an approach to de-identify medical free text using LLMs. The authors benchmarked eight locally deployable LLMs, including Llama-3 8B, Llama-3 70B, Llama-2 7B, Llama-2 70B, and Mistral 7B, on a dataset of 100 clinical letters from a German hospital. They developed the LLM-Anonymizer pipeline, which achieved a success rate of 98.05% in removing personal identifying information using Llama-3 70B. The tool is open-source, operates on local hardware, and does not require programming skills, making it accessible and practical for use in medical institutions. The study demonstrates the potential of LLMs to effectively de-identify medical texts, outperforming traditional NLP methods and providing a robust solution for privacy-preserving data sharing in healthcare.
@@
Despite advancements in summarization and paraphrasing, some challenges persist, particularly in preserving factual accuracy and precision. Jeblick et al.[177] explored the effectiveness of using ChatGPT (version December 15th, 2022) to simplify radiology reports into language understandable by non-experts. A radiologist created three hypothetical radiology reports, which were then simplified by prompting ChatGPT. Fifteen radiologists evaluated the quality of these simplified reports based on criteria such as factual correctness, completeness, and potential harm to patients. The study used Likert scale analysis and inductive free-text categorization to assess the simplified reports. Overall, the radiologists found the simplified reports to be factually correct and complete, with minimal potential for harm. However, some issues were noted, including incorrect information, omissions of relevant medical data, and occasionally misleading or vague statements. These issues highlight the need for careful supervision by medical professionals when using language models to simplify complex medical texts. A recent study by Landman et al.[173] discusses a challenge organized by Pfizer to explore the use of LLMs for automating the summarization of safety tables in clinical study reports. Various teams employed GPT models with prompt engineering techniques to generate summary texts. The datasets included safety outputs from 72 reports from recent clinical studies, split into 70% for training and 30% for testing. The study concluded that while LLMs show promise in automating the summarization of clinical study report tables, human involvement and further research are necessary to optimize their application.
@@ 3.6 Conversation
The task of conversation, or dialogue generation, can be formulated as follows. Given a dialogue history $H=(h_{1},h_{2},\ldots,h_{n})$, where each $h_{i}$ represents an utterance in the conversation, the objective is to generate an appropriate response $R$. This can be expressed as:
$R=\operatorname{argmax}_{r}P(r\mid H),$
Where $P(r\mid H)$ represents the conditional probability of generating the response $r$ given the dialogue history $H$. Typically, pre-trained decoder-based large language models are fine-tuned using specialized datasets to develop their conversational capabilities. In the medical domain, conversational applications facilitate interactive communication with patients. For example, conversational AI can be deployed as virtual health assistants that provide initial consultations based on symptoms described by patients. These systems can ask relevant follow-up questions, assess symptoms, and offer preliminary advice or direct patients to seek professional care when necessary. Additionally, these conversational tools can be utilized for patient education, explaining complex medical conditions and treatments in simple language to enhance understanding and compliance. Another notable application is in mental health support, where conversational AI can offer coping strategies and basic support, thereby augmenting traditional therapy sessions.
@@ 3.6.1 Chatbots and Health Assistants
Cung et al.[185] assessed the performance of three commercial systems - ChatGPT, Bing (Copilot), and Bard (Gemini) in the context of skeletal biology and disorders. The study involved posing 30 questions across three categories, with the responses graded for accuracy by four reviewers. While ChatGPT 4.0 had the highest overall median score, the study revealed that the quality and relevance of responses from all three chatbots varied widely, presenting issues such as inconsistency and failure to account for patient demographics. Another study explored using ChatGPT for patient-provider communication. A survey of 430 participants found that ChatGPT responses were often indistinguishable from those of healthcare providers, indicating a level of trust in chatbots for answering lower-risk health questions [186].
@@
Another research direction in the field of medical chatbots is the implementation of conversational agents specifically designed for the medical domain. Abbasian et al.[135] proposed a complex LLM-based multimodal framework for CHAs, concentrating on critical thinking, knowledge acquisition, and multi-step problem-solving. This framework aims to enable CHAs to provide personalized healthcare responses and handle intricate tasks such as stress level estimation.
Domain-specific LLMs, such as the ChatDoctor model [116], can also serve as chatbots. This model integrates a self-directed information retrieval mechanism, allowing it to access up-to-date information from online and curated offline medical databases. Evaluated using BERTScore[218], ChatDoctor exhibited a higher F1 score compared to ChatGPT-3.5, demonstrating the effectiveness of smaller domain-specific models as alternatives to large commercial solutions.
@@ 3.6.2 Mental Health Bots
The idea of using a machine as a personal psychologist dates back to at least the 1960s when Weizenbaum proposed a simple rule-based system called ELIZA [220]. Contemporary advancements in mental health chatbots are largely driven by LLMs. Yang et al.[183] investigated the capabilities of current LLMs in automated mental health analysis. Their study involved evaluating LLMs across diverse datasets for tasks such as emotional reasoning and detecting mental health conditions, employing various similarity metrics including BLEU, ROUGE family, BERTScore derivatives, BART-score [221], and human assessments. They discovered that while ChatGPT displays robust in-context learning abilities, it still encounters challenges in emotion-related tasks and requires careful prompt engineering to improve performance to enhance its performance.
@@
Saha et al. [182] introduced a Virtual Assistant for supporting individuals with Major Depressive Disorder, using a dataset called MotiVAte. Their system, based on modified GPT-2 model and reinforced learning, shows promising results in generating empathetic and motivational responses, as evidenced by both automated evaluations based on text similarity and human evaluations based on fluency, adaptability, and degree of motivation. Sharma et al.[222] introduced a dataset for training a GPT-3-based model for generating reframes with controlled linguistic attributes. Deployed on the Mental Health America website [^6] https://mhanational.org/, this allowed for a randomized field study to gather findings on human preferences. Another team explored the fine-tuning of open-source LLMs on psychotherapy assistant instructions, using a dataset from Alexander Street Press [^7] https://alexanderstreet.com/ therapy and counseling sessions. Their results indicated that LLMs fine-tuned on domain-specific instructions surpassed their non-fine-tuned counterparts in psychotherapy tasks, underscoring the significance of professional and context-specific training for these models [184].
@@
Promising outcomes have been observed through collaborations between humans and AI. A recent study[223] conducted a randomized controlled trial involving human peer supporters, demonstrating that an AI-in-the-loop agent led to a 19.60% increase in conversational empathy in interactions between individuals seeking mental health support and support specialists. This was achieved by providing suggestions for response improvements to peer supporters. This research reveals that human-AI collaboration is a crucial area for potential exploration, particularly in the medical domain.
@@
The evolving reasoning capabilities of LLMs have sparked interest in their use for disease diagnostics. Levine et al.[187] conducted experiments with the GPT-3 model to assess its diagnostic and triage accuracy. Their results indicate that GPT-3’s diagnostic accuracy is comparable to that of physicians but lags in triage accuracy. GPT-3 correctly identified the diagnosis in its top three choices for 88% of the cases, surpassing non-experts (54%) but slightly underperforming compared to professional physicians (96%). In triage performance, GPT-3 achieved an accuracy of 70%, on par with non-experts (74%) but significantly lower than physicians (91%). Despite GPT-3’s notable performance, the study raises ethical concerns, particularly regarding the model’s potential to perpetuate existing data biases, exhibiting racial and gender biases and occasionally producing misleading or incorrect information.
@@
A recent study by Liu et al. [224] introduced a framework named PharmacyGPT, which leverages the current GPT family models to emulate the role of clinical pharmacists. This research utilized real data from the ICU at the University of North Carolina Chapel Hill (UNC) Hospital. PharmacyGPT was applied to tackle various challenges in the realm of pharmacy, encompassing patient outcome studies, AI-based medication prescription generation, and interpretable patient clustering analysis. The study revealed that the GPT-4 model, when provided with dynamic context and similar samples, attained the highest accuracy among all models tested. However, the precision and recall scores were not notably high across the approaches. This outcome may be caused by the binary nature of mortality prediction, a significant imbalance in the dataset, and the complex, individualized nature of ICU pharmacy regimens. The research highlights the need for custom evaluation metrics to assess the performance of AI-generated medication plans, enhancing understanding of the models’ strengths and limitations.
4 Discussion